OpenStudy (anonymous):

How do you write the square root of 128 as a decimal rounded to the hundreth

OpenStudy (jdoe0001):

\(\bf \sqrt{128} \approx 11.{\color{red}{ 31}}37085\)

OpenStudy (jdoe0001):

100 th <-- 2 zeros, 2 decimal places, rounded up

OpenStudy (anonymous):

How about the square root of 128 as a simplified radical, no decimals.

OpenStudy (jdoe0001):

well... one can say that \(\bf \sqrt{128} \implies 2\cdot 2\cdot 2\cdot 2\cdot 2\cdot 2\cdot 2\implies 2^2\cdot 2^2\cdot 2^2\cdot 2\\ \quad \\ \implies (2^2)^3\cdot 2\implies \\ \quad \\ \sqrt{128}\implies \sqrt{(2^3)^2\cdot 2}\implies \large \sqrt[{\color{red}{ 2}}]{(2^3)^{\color{red}{ 2}}\cdot 2}\)

OpenStudy (jdoe0001):

what do you think we can take out of the radical?